Grade 12th passAlgebra

Verify that: -(-x) = x for.
x = -13/17
Verify that: -(-x) = x for.
x = -13/17

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

-13/17
We have, x = -13/17
The additive inverse of x is–x (asx+(-x)=0)
Then, the additive inverse of -13/17 is 13/17 (as 11/15 + (-11/15) = 0)
The same equality (-13/17 + 13/17) = 0, shows that the additive inverse of 13/17 is -13/17.
Or, - (13/17) = -13/17,
i.e., -(-x) = x

Thanks